OpenStudy (anonymous):

Any quantity raised to the power of zero is equal to _____. A. 0 B. 1 C. –1 D. ∞ (infinite)

OpenStudy (inkyvoyd):

let's look at it this way \(2^1=2;2^2=4;2^3=8;2^4=16\) so \(2^5=2^4*2=32\) what if you went backwards? \(\Huge 2^3=8;2^2=2^3/2=4\) \(\Huge 2^1=2^2/2=2\) so what happens if you go back one more? \(\Huge 2^0=?\)

OpenStudy (anonymous):

So is it B?

OpenStudy (inkyvoyd):

basically.

OpenStudy (inkyvoyd):

this question is actually badly phrased, because \(0^0\) is usually considered undefined or indeterminant, though advanced mathematics (calculus one sided limits) sometimes assigns it the value of 1.

OpenStudy (squirrels):

Example: \[2^{3}=3\times3\times3\] \[2^0=\] Ya feel?

OpenStudy (inkyvoyd):

Squirrels, I once saw one of your kind on my doorstep. I should've shot him with my airsoft gun when I had the chance.

OpenStudy (squirrels):

I'll kill you

OpenStudy (inkyvoyd):

I'm a hedgehog, I always win fights against puny rodents

OpenStudy (squirrels):

Not this time

OpenStudy (inkyvoyd):

:o

OpenStudy (anonymous):

Not really, but I kinda get the idea I guess. thanks!! @Squirrels

OpenStudy (anonymous):

And @inkyvoyd, thank u

OpenStudy (squirrels):

Np ↨

OpenStudy (inkyvoyd):

you're welcome @Annakat0

OpenStudy (anonymous):

@Squirrels u should fan me, so we can chat sometime yeah?

OpenStudy (squirrels):

I have diabetes

OpenStudy (anonymous):

Lolol it's alright, so do I ;b

OpenStudy (squirrels):